### v1

# jina-clip-v2: Multilingual Multimodal Embeddings for Text and Images

Abstract
Contrastive Language-Image Pretraining (CLIP) is a highly effective method for aligning images and texts in a shared embedding space. These models are widely used for tasks such as cross-modal information retrieval and multi-modal understanding. However, CLIP models often struggle with text-only tasks, underperforming compared to specialized text models. This performance disparity forces retrieval systems to rely on separate models for text-only and multi-modal tasks. In this work, we build upon our previous model, jina-clip-v1, by introducing a refined framework that utilizes multi-task, multi-stage contrastive learning across multiple languages, coupled with an improved training recipe to enhance text-only retrieval. The resulting model, jina-clip-v2, outperforms its predecessor on text-only and multimodal tasks, while adding multilingual support, better understanding of complex visual documents and efficiency gains thanks to Matryoshka Representation Learning and vector truncation. The model performs comparably to the state-of-the-art in both multilingual-multimodal and multilingual text retrieval benchmarks, addressing the challenge of unifying text-only and multi-modal retrieval systems.
1 Introduction
Contrastive text-image pre-training has become a widely adopted approach to building robust text-image alignment models. These models are particularly effective for tasks like cross-modal retrieval and zero-shot classification (Radford et al., 2021), demonstrating strong generalization capabilities. Conventional text-image pre-training excels at aligning text and image embeddings but struggles with aligning text-to-text embeddings (Koukounas et al., 2024). This is because cross-modal retrieval training is inadequate for text-only retrieval. The texts available are typically image captions, which are generally short and information-poor, and training procedures don’t use standard techniques for training text embeddings, like hard negatives. (Zhang et al., 2024a)
To overcome these limitations, Koukounas et al. (2024) introduced a multi-task, multi-stage contrastive learning approach to simultaneously align text-text and text-image embeddings. Their method involves three stages: optimizing text-text and text-image embeddings for short image-caption and text-text pairs (stage 1), refining alignment using long text-text pairs and detailed image-caption pairs (stage 2) and further enhancing performance with text triplets containing hard negatives and detailed image-caption pairs (stage 3). The resulting model, jina-clip-v1, achieves strong performance on both the cross-modal CLIP Benchmark111https://github.com/LAION-AI/CLIP_benchmark and text embedding MTEB Benchmark (Muennighoff et al., 2023).
Despite its strengths, this model has several limitations. First, it is an English-only multimodal embedding model, which makes it unsuitable for multilingual document retrieval. Second, jina-clip-v1 struggles with images that contain text, tables, graphs, diagrams, and other complex content (Faysse et al., 2024).
To address these challenges, we propose an enhanced text-image pre-training scheme. Our approach incorporates multilingual text-image and text-text pairs, along with text-image pairs containing text and other difficult-to-process complex visual structures. Additionally, we gradually increase image resolution during pre-training to balance computational efficiency and model performance. The resulting model, jina-clip-v2 depicted in Figure 1, not only achieves competitive performance on cross-modal retrieval benchmarks against state-of-the-art models like NLLB-CLIP-SigLIP (Visheratin, 2023b) but also performs comparably to dedicated multilingual text embedding models like jina-embeddings-v3 (Sturua et al., 2024) on the Retrieval and STS tasks of the multilingual MTEB benchmark (Muennighoff et al., 2023). Moreover, due to the inclusion of visually rich training data and the progressive increase in image resolution, jina-clip-v2 demonstrates significantly improved performance on benchmarks for visually rich document retrieval, compared to jina-clip-v1. To reduce vector storage costs, we employ Matryoshka Representation Learning (Kusupati et al., 2024) to enable dimension truncation on the output vectors with minimal performance degradation.
2 Related Work
Radford et al. (2021) pioneered the dual encoder architecture trained contrastively on image-text pairs, demonstrating impressive zero-shot performance and laying the groundwork for pre-trained vision-language models. Building upon this foundation, Zhai et al. (2023) propose an alternative and more efficient sigmoid objective, while Sun et al. (2023) optimize the model’s dimensions and scale up both dataset and model sizes, achieving state-of-the-art performance on cross-modal tasks. Koukounas et al. (2024) enhance a CLIP-based retriever with strong text-to-text and cross-modal retrieval capabilities, though fall short in addressing retrieval scenarios that involve candidate pools with heterogeneous modalities, due to the modality gap (Liang et al., 2022). To mitigate the modality bias, Lin et al. (2024) take a different approach by fine-tuning a multimodal LLM with modality-aware hard-negative mining, creating a universal multimodal retriever and maintaining competitive text-to-text retrieval performance across diverse tasks.
Extensive research has also focused on extending CLIP’s capabilities to multilingual contexts. To address the lack of image-caption pairs in languages other than English, Carlsson et al. (2022) apply knowledge distillation to retrain the text encoder using machine translation data. NLLB-CLIP (Visheratin, 2023a) leverages Locked-image Tuning (LiT) (Zhai et al., 2022) alongside the NLLB (Costa-jussà et al., 2022) text encoder, achieving state-of-the-art results in both retrieval and classification tasks.
The importance of good information retrieval in Retrieval-Augmented Generation (RAG) systems has motivated rapid progress in multilingual text embedding models. These models are typically based on either a robust encoder-only architecture, such as XLM-RoBERTa (Conneau et al., 2020) or a decoder-only multilingual large language model, such as Mistral 7B (Jiang et al., 2023). Multilingual E5 (Wang et al., 2024) and BGE-M3 (Chen et al., 2024) both use XLM-RoBERTa as their backbone, leveraging extensive multilingual training with instruction tuning and multi-task learning techniques, respectively. mGTE (Zhang et al., 2024b) achieves comparable performance to BGE-M3 using a smaller transformer architecture, enhanced by RoPE (Su et al., 2023) to extend the context length. Similarly, Sturua et al. (2024) employ RoPE to increase the model’s context length and apply LoRA tuning (Hu et al., 2021) to optimize embeddings for downstream tasks.
3 Training
| Feature | Text Encoder | Image Encoder |
| Base Model | Jina-XLM-RoBERTa (Sturua et al., 2024) | EVA02 L/14 (Sun et al., 2023) |
| Parameters | 561M | 304M |
| Input Specification | 8,192 tokens (max) | 512×512 pixels |
| Output Dimensions | 1,024 | 1,024 |
| Layers | 24 | 24 |
| Attention Implementation | FlashAttention2 (Dao et al., 2022) | xFormers (Lefaudeux et al., 2022) |
| Pooling Strategy | Mean pooling | CLS pooling |
| Additional Features | 89 languages supported | Patch size 14x14 |
The jina-clip-v2 model uses the dual encoder architecture, introduced in the original CLIP (Radford et al., 2021) model and reused in jina-clip-v1 (Koukounas et al., 2024). One encoder is responsible for encoding texts in a 1024-dimensional embedding space, while a second encoder embeds images in the same embedding space.
The text encoder is initialized with the pre-trained Jina-XLM-RoBERTa model weights. Introduced in jina-embeddings-v3 (Sturua et al., 2024), the Jina-XLM-RoBERTa model is a port of the XLM-RoBERTa (Conneau et al., 2020) weights to a modern encoder-only architecture that adds Flash Attention (Dao et al., 2022), rotary positional embeddings (Su et al., 2023) and LoRA (Hu et al., 2021).
For the image encoder, we have opted for the EVA02 (Fang et al., 2023b) family of ViT models, similarly to jina-clip-v1. We have selected a relatively large pre-trained model, similar in size to the text encoder. This model implementation includes 2D rotary positional embeddings (Su et al., 2023) and a memory-efficient attention implementation based on xFormers (Lefaudeux et al., 2022).
Architectural details for both encoders are presented in Table 1.
3.1 Datasets
Similarly to jina-clip-v1 (Koukounas et al., 2024), we construct four different training datasets. Dataset denotes a dataset of text pairs, denotes a dataset of text samples with hard negatives, a multimodal short-caption dataset and a multimodal long-caption dataset. All datasets are multi-lingual.
The text corpora and are introduced in Sturua et al. (2024) and contain data in 30 languages. The text pair corpus consists of query and target pairs from a diverse collection of text-pair datasets. The triplet text corpus is a high-quality dataset with hard negatives from various sources. Each training item contains one annotated positive and seven negative items. Regarding , we focus on Retrieval and Semantic Textual Similarity (STS) datasets for the scope of this work.
Our multi-modal training data comes from multiple sources. For the English part, we rely on the DFN dataset (Fang et al., 2023a), a collection of 400M image-caption pairs, filtered using data filtering networks. For the rest of the languages, we collect 400M samples from CommonPool (Gadre et al., 2023), using multilingual CLIP to identify and discard low-quality items. All images are resized to . These two large-scale datasets are included in and a small part with captions longer than 256 tokens is held out for .
Inspired by visual QA applications (Faysse et al., 2024), We have diversified our training data with PDFs, scientific graphs, infographics and Wikipedia images. The following datasets are included in both and : DocVQA (Mathew et al., 2021b), TatDQA (Zhu et al., 2022), InfographicsVQA (Mathew et al., 2021a), SciGraphQA (Li & Tajbakhsh, 2023), ArXivQA and ArXivCAP (Li et al., 2024), WIT (Srinivasan et al., 2021) and ViDoRe synthetic training data (Faysse et al., 2024). For the QA datasets, to obtain a caption for an image, we concatenate query and answer for each sample. For the WIT dataset, we use the reference caption in and a concatenation of caption, page title, section tile, page description and section description for .
Finally, includes multilingual synthetic long captions, similar to how Koukounas et al. (2024) make use of ShareGPT4v (Chen et al., 2023a). We use the GPT4v API (OpenAI, 2023) to generate detailed image descriptions for 40,000 images in 30 languages. This adds 1.2M generated multilingual long captions to .
3.2 Training Stages
Following the strategy proposed in Koukounas et al. (2024), we employ a multi-task, multi-stage training approach, in order to optimize the model for two tasks simultaneously: text-image matching and text-text matching.
Stage 1 focuses on aligning the multi-modal representations while also improving the text representations of the text encoder. We train on and , with a small context length of 77 and an image resolution of to enable large batch sizes. Towards the end, when performance peaks, we switch to a higher image resolution of using positional embedding interpolation as a warm-up for the next stage.
Stage 2 trains on and , with context length increased from 77 to 512 and image resolution set at . This stage improves text embedding performance on longer text lengths, while maintaining alignment between the two modalities.
Stage 3 uses hard negatives from to further improve the text encoder in distinguishing relevant from irrelevant text. To maintain text-image alignment, we continue training with . We do one more positional embedding interpolation to increase the image resolution to and train on this resolution for the duration of stage 3.
Appendix table 6 specifies the training details for each step.
3.3 Loss Functions
Given a batch of embedding pairs , the InfoNCE loss (Van den Oord et al., 2018), given in equation 1, evaluates the cosine similarity between query and its corresponding target , relative to the similarity of all other targets in the batch. The loss is calculated in both directions to preserve the symmetry of similarity measures. The temperature parameter influences how the loss function weighs minor differences in the similarity scores.
When hard negatives are available for query , an extended version of the loss is used. Given a batch of samples , consisting of a query , a positive match , and seven negatives , the extended loss function, denoted here as , is given in equation 2. Similarly to , this loss function is bidirectional.
Equation 3 formulates the loss for each of the three stages. In each stage, we optimize a joint loss function, i.e. the sum of two loss functions, one operating on text representations and one on multi-modal representations. All stages optimize except Stage 3, which uses hard negatives in the text loss branch and thus calculates .
In Equation 3, denotes a batch of text pairs drawn from the dataset of text pairs , while denotes a batch of text samples with negatives drawn from the hard negatives dataset . Similarly denotes a batch of text-image pairs drawn from the multimodal short-caption dataset , while denotes a batch of text-image pairs sampled from the multimodal long-caption dataset .
3.4 Matryoshka Representation Learning
In all stages of training, each loss function is evaluated on the full 1024-dimension output vectors of the two encoders as well as on truncated vectors of 64, 128, 256, 512 and 768 dimensions. In other words, each loss component on equation 3 is the sum of the same loss function evaluated at different embedding dimensionalities.
This training technique, called Matryoshka Representation Learning, (Kusupati et al., 2024) trains the model to learn the most important features in the vector sequence. This makes it possible, at inference time, to truncate embedding vectors, thereby reducing storage and computation costs when comparing vectors, at a minor performance penalty.
4 Evaluation
We have evaluated the performance of jina-clip-v2 on a selection of benchmarks: English and multilingual cross-modal retrieval benchmarks in Section 4.1, common text embedding tasks from the MTEB benchmark suite in Section 4.2, visual document retrieval tasks in Section 4.3, and an ablation study of jina-clip-v2’s matryoshka representations in Section 4.4.
Detailed evaluation results are presented in the appendices. Detailed CLIP Benchmark results are given in Appendix tables 7, 8 and 9, MTEB results are in in Appendix 10. ViDoRe benchmark results are in Appendix 11. A full MRL evaluation can be found in tables 12, 13, 14, 15 in the appendices.
4.1 Cross-modal Evaluation
| Benchmark | CLIP Benchmark - EN | Crossmodal-3600 | Multilingual MS COCO | |||
| Task Type | Zero-Shot Retrieval | |||||
| Model - Metric | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 |
| nllb-siglip-large | 81.54 | 88.15 | 82.07 | 80.16 | 87.60 | 85.37 |
| nllb-siglip-base | 79.57 | 86.38 | 79.29 | 76.56 | 86.23 | 84.87 |
| jina-clip-v1 | 77.75 | 87.65 | - | - | - | - |
| jina-clip-v2stage 1 | 73.87 | 86.61 | 73.51 | 79.64 | 80.66 | 83.02 |
| jina-clip-v2stage 2 | 79.81 | 89.57 | 84.13 | 84.12 | 86.11 | 86.45 |
| jina-clip-v2 | 79.09 | 89.73 | 81.43 | 83.23 | 84.87 | 86.03 |
T-I r@5 : Text to Image Recall@5 [%] I-T r@5 : Image to Text Recall@5 [%]
We evaluate jina-clip-v2 on a set of English and multilingual cross-modal tasks, conducting a comparative analysis against both its predecessor jina-clip-v1 and the NLLB-CLIP-SigLIP (Visheratin, 2023b) large and base variants, a top-performing multilingual CLIP model. Additionally, we track the model’s performance at different training stages to evaluate the effectiveness of different parts of the training protocol. Measuring 865M parameters, our model sits between the two NLLB-CLIP-SigLIP versions in terms of size: nllb-siglip-base (507M parameters, 41% smaller than ours) and nllb-siglip-large (1.2B parameters, 39% larger than ours).
For English zero-shot image-text and text-image retrieval, we conduct evaluations on Flickr30K (Young et al., 2014) and MS COCO Captions (Chen et al., 2015), both of which are included in the CLIP Benchmark. For multilingual cross-modal tasks, we assess performance on the Crossmodal-3600 (Thapliyal et al., 2022) and multilingual MS COCO (Chen et al., 2015) datasets.
As shown in Table 2, jina-clip-v2 demonstrates strong performance across both English and multilingual benchmarks while showing clear progression through its training stages. On the English CLIP Benchmark, the final model achieves 79.09% and 89.73% on text-to-image and image-to-text retrieval respectively, showing significant improvement from stage 1’s 73.87% and 86.61%. While nllb-siglip-large leads in text-to-image retrieval at 81.54%, jina-clip-v2 outperforms its predecessor jina-clip-v1 and surpasses both NLLB-CLIP-SigLIP variants in image-to-text retrieval by 2.1% and 1.6% respectively. On the multilingual front, we obtain competitive results on Crossmodal-3600, achieving 81.43% and 83.23% on text-to-image and image-to-text retrieval respectively, approaching the performance of nllb-siglip-large. On multilingual MS COCO (Chen et al., 2015), jina-clip-v2 performs comparably to NLLB-CLIP-SigLIP models, reaching 84.87% and 86.03% on text-to-image and image-to-text retrieval. The improvement from Stage 1 to the final model is notable across all benchmarks, with gains of up to 10.6 percentage points on Crossmodal-3600’s text-to-image retrieval, demonstrating the effectiveness of the training strategy.
4.2 Text Retrieval and STS Evaluation
| Benchmark | Retrieval | STS | ||
| Metric | nDCG@10 [%] | Spearman Correlation [%] | ||
| Model - Language | EN | Multilingual | EN | Multilingual |
| nllb-siglip-large | 24.91 | - | 74.89 | - |
| jina-embeddings-v3 | 53.87 | 72.58 | 85.80 | 69.81 |
| jina-clip-v1 | 48.33 | - | 80.92 | - |
| jina-clip-v2stage 1 | 40.91 | - | 79.67 | - |
| jina-clip-v2stage 2 | 43.17 | - | 80.33 | - |
| jina-clip-v2 | 49.32 | 69.85 | 81.29 | 67.77 |
Table 3 presents the evaluation on MTEB tasks, demonstrating the effectiveness of jina-clip-v2’s text encoder for both retrieval and semantic textual similarity (STS) tasks. On English retrieval tasks (nDCG@10), jina-clip-v2 achieves an average score of 49.32, a significant improvement over nllb-siglip-large’s 24.91 and a slight improvement over jina-clip-v1’s 48.33, though not reaching jina-embeddings-v3’s score of 53.87. For multilingual retrieval, the model achieves 69.85, approaching jina-embeddings-v3’s performance of 72.58. In STS tasks measured by Spearman Correlation, jina-clip-v2 demonstrates competitive performance with 81.29 for English tasks, surpassing nllb-siglip-large’s 74.89 and comparable to jina-clip-v1’s 80.92, though below jina-embeddings-v3’s 85.8. On multilingual STS tasks, the model achieves 67.77, slightly below jina-embeddings-v3’s 69.8.
4.3 Visual Document Retrieval
| Model | Average (nDCG@5) [%] |
| jina-clip-v1 | 17.72 |
| nllb-siglip-large | 46.55 |
| jina-clip-v2 | 52.65 |
| jina-clip-v2stage 1 | 37.37 |
| jina-clip-v2stage 2 | 47.76 |
Visual document retrieval requires vision models to capture fine-grained information from images to accurately match documents with queries. This task is particularly challenging as it deals with high-resolution document images, which differ significantly from typical image-caption datasets. As shown in table 4, jina-clip-v2 achieves state-of-the-art performance on the ViDoRe Benchmark (Faysse et al., 2024) for visual document understanding, with an average nDCG@5 score of 52.65%, significantly outperforming both its predecessor jina-clip-v1 (17.72%) and nllb-siglip-large (46.55%). The progression through training stages demonstrates substantial improvements, advancing from 37.37% in stage 1 to 47.76% in stage 2, before reaching peak performance in the final model. These results highlight the effectiveness of our training strategy and the impact of the new datasets introduced for document retrieval tasks.
4.4 Matryoshka Representation Learning
| Dataset - Dimension |
| Text-Image Retrieval - Recall@5 [%] | ||||||||||||
| Avg CLIP Benchmark | 79.10 | 79.12 | 78.93 | 78.32 | 75.90 | 70.51 | ||||||
| Avg Crossmodal-3600 | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
| Avg Multilingual MS COCO | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
| Image-Text Retrieval - Recall@5 [%] | ||||||||||||
| Avg CLIP Benchmark | 89.73 | 89.60 | 89.55 | 89.35 | 87.48 | 83.20 | ||||||
| Avg Crossmodal-3600 | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
| Avg Multilingual MS COCO | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
| Text-Text Retrieval - nDCG@10 [%] | ||||||||||||
| Avg Retrieval | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
| Semantic Textual Similarity - Spearman Correlation [%] | ||||||||||||
| Avg STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
Table 5 presents the impact of embedding truncation on the performance of our model. The table is divided into 4 task categories: text-to-image retrieval, image-to-text retrieval, text-to-text retrieval and semantic textual similarity (STS).
The results show that performance remains highly stable when reducing dimensions from 1024 to 256, with minimal degradation (typically less than 1%) across all evaluation tasks. At 256 dimensions, representing a 75% reduction in embedding size, the model effectively preserves essential semantic information. Substantial performance degradation is only observed at very low dimensions (128 and 64), indicating the model’s very robust acquisition of image and text features.
5 Analysis
5.1 The Role of Image Resolution in Complex Document Retrieval
Our empirical study demonstrates that image resolution plays a critical role in retrieving visually rich documents. To better understand the relationship between image resolution and model performance on such documents, we have conducted a targeted experiment.
We checkpointed jina-clip-v2 at the end of Stage 1, when it had only seen images with a resolution of (224, 224). This model checkpoint was trained on 1.45 billion image-caption pairs and 1.45 billion text-text pairs. We conducted four runs, keeping the image resolution to for the first run and increasing to , , and for three additional runs. Each run is trained for 3500 steps using the same visually rich training set and hyperparameters. The models are then evaluated on the ViDoRe benchmark (Faysse et al., 2024), which comprises 10 datasets designed for retrieving visually rich documents, based on text queries.
We have plotted the results in Figure 2. Unsurprisingly, increasing image resolution has a positive impact on linking captions to visually rich documents. The most significant improvement occurs when resolution increases from to , with the average nDCG@5 score across 10 benchmarks rising from to . A further increase in resolution to also brings a noticeable gain in performance.
Considering that the model is trained with a patch size of 14, increasing the resolution beyond significantly raises the number of patches, a increase when moving from to . This leads to substantial additional computational overhead while yielding only marginal improvement, with nDCG@5 increasing by just . Therefore, we identify as an optimal image resolution, offering a good balance between performance and computational efficiency.
5.2 Unified Batch vs Multi-Task Learning
Our training approach, as outlined in Section 3, involves two tasks at each stage, both optimized in a multi-task learning framework: a multi-modal alignment task for images and text, and second text-only task. Both use the InfoNCE (Van den Oord et al., 2018) objective. Regardless of whether the loss is computed on image-text or text-text pairs, the same loss function is applied to pairs of vectors. This raises the question: Would unifying the two batches of pairs into a single batch, irrespective of modality, and optimizing a single contrastive objective, offer any benefits?
We refer to this approach as the unified batch technique, in contrast to the multi-task learning paradigm. We present the training paradigm in Figure 3. The motivation is two-fold: first, simplicity of computation and secondly, modality gap mitigation (Liang et al., 2022) by using in-batch negatives across modalities. Compared to multi-task training, the unified batch approach utilizes a single InfoNCE loss with a shared temperature value, attempting to force all modalities to align within the same embedding space. Additionally, we integrate a SimCLR-like self-supervised learning method (Chen et al., 2020) into the contrastive training, aiming to provide more robust image representation.
However, our Stage 1 experiments show that this technique does bring some early improvements, but fails beyond a certain point, at least with regard to cross-modal tasks. This was reflected in the slow decay of the loss temperature (our temperature is a trainable parameter in this case) and its eventual stabilization around . This is in contrast to the multi-task learning approach, where the temperature value decreases more rapidly and keeps decreasing throughout training, leading to better performance on cross-modal tasks.
We hypothesize that this limitation arises from the fundamental information asymmetry between the visual and textual modalities (Schrodi et al., 2024). Images include fine-grained details of information whereas the corresponding textual descriptions include more coarse-grained level of information (Schrodi et al., 2024; Chen et al., 2023b). In temperature-scaled contrastive learning, the temperature parameter controls the hardness of the loss function with respect to the negative samples, i.e., the level of penalties on the hard negative samples (Wang & Liu, 2021). From an information-theory perspective, the level of penalty on the hard negative text-text samples theoretically differs from that of image-text samples due to the information gap problem. Consequently, as supported by our experiments, enforcing a unified temperature parameter across both modalities is potentially theoretically suboptimal.
6 Conclusion
In this work we present an enhanced strategy for training dual-encoder vision-language embedding models for cross-modal and text retrieval tasks, using contrastive learning. We introduce improvements to the pipeline, namely support for multiple languages, Matryoshka Representation Learning, and finer visual perception. We have trained and released a new model, jina-clip-v2, which has strong cross-modal and text retrieval performance on standard benchmarks. Finally, we have conducted analyses of three important considerations for CLIP models going forward: the ability to understand complex visual inputs performance how the image resolution affects it, and the effects of efforts to mitigate the modality gap.
Appendix A Appendix
| Parameter | Stage 1 | Stage 2 | Stage 3 |
| Image encoder weights init | EVA02 ViT L/14 (Sun et al., 2023) | Stage 1 | Stage 2 |
| Text encoder weights init | JinaXLMRoBERTa (Sturua et al., 2024) | Stage 1 | Stage 2 |
| Peak image encoder LR | 2e-4 | 5e-5 | 5e-6 |
| Image encoder layer-wise LR decay | 1 | 0.98 | 1 |
| Peak text encoder LR | 1e-4 | 5e-5 | 1e-4 |
| Text encoder layer-wise LR decay | 1 | 0.98 | 1 |
| Image-text pairs batch size | |||
| Text pairs batch size | |||
| Total steps | |||
| Max sequence length | |||
| Image-text pairs samples seen | 1.7B | 50M | 16M |
| Text pairs samples seen | 1.7B | 50M | 2M |
| Number of GPUs - H100s 80GB | 8 | 8 | 8 |
| LR schedule | cosine decay | ||
| Optimizer | AdamW (Loshchilov & Hutter, 2019) | ||
| Optimizer hyper-parameters | |||
| Weight decay | 0.02 | ||
| Input resolution | (224…384, 224…384) | (384, 384) | (512, 512) |
| Patch size | (14, 14) | ||
| Numerical precision | bfloat16 | ||
| Dataset - Model |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||||
| Flickr30K (Young et al., 2014) | 89.84 | 86.84 | 90.04 | 89.04 | 92.24 | 90.02 | ||||||||||
| MS COCO (Chen et al., 2015) | 68.35 | 60.91 | 69.59 | 66.42 | 70.84 | 69.13 | ||||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||||
| Flickr30K (Young et al., 2014) | 98.00 | 96.10 | 97.40 | 96.40 | 97.10 | 95.00 | ||||||||||
| MS COCO (Chen et al., 2015) | 81.46 | 77.12 | 81.74 | 79.02 | 79.20 | 77.76 | ||||||||||
| Language - Model |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||
| average | 81.43 | 73.51 | 84.13 | 82.07 | 79.29 | |||||||||
| ar | 73.56 | 66.22 | 76.89 | 78.92 | 76.94 | |||||||||
| bn | 63.78 | 50.19 | 68.11 | 75.19 | 74.19 | |||||||||
| da | 85.39 | 76.53 | 87.97 | 87.14 | 86.64 | |||||||||
| de | 91.25 | 84.64 | 92.42 | 89.56 | 87.25 | |||||||||
| el | 75.03 | 68.53 | 77.92 | 77.83 | 71.97 | |||||||||
| en | 75.83 | 69.78 | 77.78 | 73.11 | 72.22 | |||||||||
| es | 83.64 | 78.28 | 86.28 | 82.64 | 79.97 | |||||||||
| fi | 82.83 | 75.28 | 85.89 | 86.42 | 81.44 | |||||||||
| fr | 88.78 | 83.50 | 90.67 | 87.86 | 85.58 | |||||||||
| hi | 55.25 | 42.03 | 59.47 | 60.31 | 58.08 | |||||||||
| id | 84.22 | 73.69 | 86.94 | 86.31 | 84.17 | |||||||||
| it | 88.33 | 81.19 | 89.64 | 85.94 | 82.67 | |||||||||
| ja | 87.03 | 77.28 | 90.03 | 86.06 | 83.14 | |||||||||
| ko | 78.81 | 71.81 | 83.22 | 78.75 | 75.47 | |||||||||
| nl | 82.56 | 75.72 | 84.47 | 81.69 | 78.86 | |||||||||
| no | 81.08 | 71.97 | 83.08 | 82.69 | 79.97 | |||||||||
| pl | 84.00 | 79.33 | 86.50 | 82.72 | 78.61 | |||||||||
| pt | 82.42 | 76.17 | 85.19 | 82.69 | 79.44 | |||||||||
| ro | 89.36 | 82.92 | 92.22 | 90.03 | 86.17 | |||||||||
| ru | 88.97 | 82.97 | 91.11 | 86.44 | 83.78 | |||||||||
| sv | 78.06 | 71.33 | 80.56 | 79.33 | 76.17 | |||||||||
| th | 81.61 | 70.92 | 85.08 | 81.14 | 78.83 | |||||||||
| tr | 81.31 | 75.31 | 84.53 | 83.47 | 81.00 | |||||||||
| uk | 88.56 | 82.28 | 89.89 | 85.44 | 81.89 | |||||||||
| vi | 86.64 | 76.31 | 89.06 | 85.56 | 82.56 | |||||||||
| zh | 78.97 | 66.97 | 82.50 | 76.56 | 74.64 | |||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||
| average | 83.23 | 79.64 | 84.12 | 80.16 | 76.56 | |||||||||
| ar | 76.25 | 72.14 | 76.67 | 75.86 | 73.69 | |||||||||
| bn | 69.00 | 63.78 | 70.22 | 75.58 | 73.61 | |||||||||
| da | 88.53 | 84.42 | 88.86 | 86.53 | 84.69 | |||||||||
| de | 92.47 | 88.42 | 92.47 | 87.50 | 84.44 | |||||||||
| el | 73.33 | 73.22 | 75.61 | 74.81 | 68.89 | |||||||||
| en | 78.58 | 74.33 | 79.61 | 70.81 | 69.08 | |||||||||
| es | 86.28 | 81.19 | 87.36 | 81.19 | 77.00 | |||||||||
| fi | 84.19 | 81.17 | 85.64 | 84.25 | 78.75 | |||||||||
| fr | 90.89 | 86.56 | 91.00 | 86.64 | 83.64 | |||||||||
| hi | 61.64 | 54.39 | 61.56 | 61.89 | 59.83 | |||||||||
| id | 86.31 | 83.64 | 87.64 | 84.33 | 81.89 | |||||||||
| it | 90.17 | 84.31 | 90.53 | 83.50 | 78.64 | |||||||||
| ja | 88.50 | 85.53 | 89.44 | 84.03 | 81.14 | |||||||||
| ko | 81.42 | 79.22 | 83.00 | 76.75 | 74.17 | |||||||||
| nl | 82.47 | 77.94 | 83.33 | 79.28 | 74.33 | |||||||||
| no | 83.75 | 77.22 | 84.42 | 82.08 | 77.03 | |||||||||
| pl | 84.61 | 83.72 | 85.69 | 79.58 | 75.39 | |||||||||
| pt | 83.94 | 79.31 | 84.06 | 79.39 | 74.89 | |||||||||
| ro | 91.31 | 87.39 | 92.08 | 88.83 | 83.58 | |||||||||
| ru | 90.64 | 87.75 | 90.89 | 84.64 | 80.61 | |||||||||
| sv | 78.28 | 77.53 | 80.11 | 76.58 | 72.31 | |||||||||
| th | 81.94 | 80.22 | 84.11 | 79.56 | 76.28 | |||||||||
| tr | 82.67 | 80.36 | 83.81 | 80.36 | 77.69 | |||||||||
| uk | 89.53 | 86.42 | 89.97 | 83.22 | 78.67 | |||||||||
| vi | 88.06 | 85.64 | 88.58 | 83.14 | 80.03 | |||||||||
| zh | 79.22 | 74.94 | 80.42 | 73.83 | 70.22 | |||||||||
| Language - Model |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||
| average | 84.87 | 80.66 | 86.11 | 87.60 | 86.23 | |||||||||
| de | 85.70 | 80.40 | 86.80 | 88.30 | 87.00 | |||||||||
| en | 89.40 | 84.00 | 89.40 | 89.40 | 88.80 | |||||||||
| es | 85.90 | 82.70 | 87.40 | 88.20 | 86.70 | |||||||||
| fr | 85.10 | 80.90 | 87.30 | 87.70 | 86.90 | |||||||||
| it | 85.80 | 83.20 | 87.10 | 89.30 | 87.80 | |||||||||
| ko | 82.10 | 78.30 | 83.00 | 85.20 | 83.60 | |||||||||
| pl | 86.50 | 81.10 | 88.00 | 89.40 | 87.60 | |||||||||
| ru | 81.10 | 79.30 | 82.10 | 83.40 | 82.40 | |||||||||
| tr | 83.70 | 81.30 | 86.30 | 88.30 | 86.80 | |||||||||
| zh | 83.40 | 75.40 | 83.70 | 86.80 | 84.70 | |||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||
| average | 86.03 | 83.02 | 86.45 | 85.37 | 84.56 | |||||||||
| de | 86.20 | 82.90 | 86.90 | 84.40 | 84.30 | |||||||||
| en | 90.50 | 87.80 | 90.60 | 88.30 | 87.50 | |||||||||
| es | 87.00 | 84.90 | 87.50 | 85.80 | 85.20 | |||||||||
| fr | 85.20 | 82.20 | 85.00 | 86.30 | 84.40 | |||||||||
| it | 88.00 | 83.10 | 87.70 | 86.70 | 85.60 | |||||||||
| ko | 82.10 | 79.30 | 84.00 | 83.20 | 82.60 | |||||||||
| pl | 88.50 | 84.50 | 88.90 | 86.80 | 86.60 | |||||||||
| ru | 83.10 | 79.60 | 81.80 | 80.90 | 81.20 | |||||||||
| tr | 85.60 | 83.20 | 85.90 | 87.10 | 85.40 | |||||||||
| zh | 84.10 | 82.70 | 86.20 | 84.20 | 82.80 | |||||||||
| Dataset - Model |
| Retrieval - nDCG@10 | ||||||||||||||||
| Average Retrieval | 49.33 | 53.87 | 40.91 | 43.17 | 24.92 | 48.33 | ||||||||||
| ArguAna | 43.55 | 54.33 | 41.87 | 45.84 | 40.63 | 49.36 | ||||||||||
| ClimateFEVER | 30.12 | 42.36 | 19.32 | 23.74 | 15.57 | 24.81 | ||||||||||
| CQADupstackRetrieval | 41.20 | 42.36 | 41.57 | 42.09 | 21.06 | 40.92 | ||||||||||
| DBPedia | 37.44 | 41.00 | 22.04 | 25.48 | 20.46 | 36.64 | ||||||||||
| FEVER | 84.96 | 89.05 | 60.97 | 66.13 | 14.98 | 76.28 | ||||||||||
| FiQA2018 | 41.93 | 47.35 | 41.87 | 40.97 | 12.90 | 38.27 | ||||||||||
| HotpotQA | 60.14 | 64.67 | 49.28 | 47.00 | 22.38 | 61.89 | ||||||||||
| MSMARCO | 37.47 | 40.82 | 23.06 | 27.05 | 13.35 | 36.91 | ||||||||||
| NFCorpus | 32.89 | 36.62 | 30.95 | 32.11 | 18.97 | 33.52 | ||||||||||
| NQ | 57.17 | 64.23 | 39.78 | 42.80 | 16.96 | 58.09 | ||||||||||
| QuoraRetrieval | 88.14 | 89.09 | 87.12 | 86.92 | 80.71 | 87.88 | ||||||||||
| SCIDOCS | 18.90 | 19.81 | 20.51 | 20.91 | 10.62 | 20.24 | ||||||||||
| SciFact | 65.34 | 72.31 | 68.26 | 67.61 | 35.28 | 67.34 | ||||||||||
| TRECCOVID | 76.73 | 77.72 | 51.29 | 58.63 | 38.15 | 71.61 | ||||||||||
| Touche2020 | 23.94 | 26.30 | 15.72 | 20.21 | 11.75 | 21.15 | ||||||||||
| STS - Spearman correlation based on cosine similarity | ||||||||||||||||
| Average STS | 81.29 | 85.80 | 79.67 | 80.33 | 74.90 | 80.92 | ||||||||||
| BIOSSES | 83.00 | 88.69 | 83.55 | 83.62 | 65.07 | 83.75 | ||||||||||
| SICK-R | 82.38 | 89.62 | 78.96 | 78.63 | 75.83 | 78.95 | ||||||||||
| STS12 | 76.72 | 82.43 | 73.41 | 74.57 | 73.60 | 73.52 | ||||||||||
| STS13 | 79.90 | 89.49 | 80.25 | 82.35 | 76.50 | 83.24 | ||||||||||
| STS14 | 77.49 | 84.95 | 75.13 | 76.89 | 76.35 | 78.68 | ||||||||||
| STS15 | 86.42 | 89.31 | 84.98 | 85.62 | 80.90 | 87.46 | ||||||||||
| STS16 | 85.18 | 86.85 | 83.70 | 82.66 | 74.73 | 83.77 | ||||||||||
| STS17 | 87.87 | 89.99 | 86.13 | 86.95 | 83.84 | 89.77 | ||||||||||
| STS22 | 67.07 | 67.28 | 66.31 | 67.58 | 62.52 | 65.15 | ||||||||||
| STSBenchmark | 86.87 | 89.44 | 84.27 | 84.40 | 79.64 | 84.93 | ||||||||||
| Task |
| Retrieval - nDCG@5 | |||||||||||||
| Average | 52.65 | 37.37 | 47.76 | 17.72 | 46.55 | ||||||||
| ArxivQ | 64.92 | 58.78 | 66.14 | 25.40 | 30.44 | ||||||||
| DocQ | 24.64 | 21.16 | 24.81 | 11.90 | 23.82 | ||||||||
| InfoQ | 57.90 | 56.62 | 56.63 | 35.50 | 59.87 | ||||||||
| TabF | 45.91 | 37.28 | 43.18 | 20.20 | 70.68 | ||||||||
| TATQ | 30.25 | 12.09 | 25.79 | 3.30 | 20.00 | ||||||||
| Shift | 34.07 | 28.41 | 31.54 | 3.80 | 30.79 | ||||||||
| AI | 68.07 | 32.95 | 55.64 | 15.20 | 47.90 | ||||||||
| Energy | 62.15 | 49.15 | 60.97 | 19.70 | 64.94 | ||||||||
| Gov | 68.97 | 37.81 | 55.56 | 21.40 | 58.62 | ||||||||
| Health | 69.05 | 39.50 | 57.37 | 20.80 | 58.43 | ||||||||
| Dataset - Dimension |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||
| Flickr30K (Young et al., 2014) | 89.84 | 89.98 | 89.70 | 89.20 | 87.22 | 81.60 | ||||||
| MS COCO (Chen et al., 2015) | 68.35 | 68.26 | 68.16 | 67.43 | 64.58 | 59.41 | ||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||
| Flickr30K (Young et al., 2014) | 98.00 | 98.10 | 98.00 | 98.00 | 96.30 | 93.40 | ||||||
| MS COCO (Chen et al., 2015) | 81.46 | 81.10 | 81.10 | 80.70 | 78.66 | 73.00 | ||||||
| Language - Dimension |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||
| average | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
| ar | 73.56 | 73.39 | 73.17 | 72.61 | 68.42 | 62.28 | ||||||
| bn | 63.78 | 63.67 | 63.64 | 62.39 | 57.58 | 49.58 | ||||||
| da | 85.39 | 85.31 | 84.67 | 84.53 | 81.69 | 75.69 | ||||||
| de | 91.25 | 91.28 | 91.47 | 91.47 | 88.75 | 84.44 | ||||||
| el | 75.03 | 75.08 | 75.25 | 74.69 | 72.00 | 66.81 | ||||||
| en | 75.83 | 75.97 | 76.03 | 75.72 | 74.03 | 69.67 | ||||||
| es | 83.64 | 83.67 | 83.86 | 83.14 | 81.00 | 76.19 | ||||||
| fi | 82.83 | 83.03 | 82.61 | 81.67 | 79.44 | 74.94 | ||||||
| fr | 88.78 | 88.75 | 88.53 | 88.28 | 86.92 | 82.14 | ||||||
| hi | 55.25 | 54.97 | 55.14 | 54.14 | 50.06 | 42.33 | ||||||
| id | 84.22 | 84.00 | 84.11 | 83.25 | 80.00 | 74.17 | ||||||
| it | 88.33 | 88.53 | 88.31 | 87.83 | 85.11 | 79.75 | ||||||
| ja | 87.03 | 87.03 | 86.89 | 86.33 | 82.92 | 75.39 | ||||||
| ko | 78.81 | 78.72 | 78.56 | 77.03 | 73.44 | 65.78 | ||||||
| nl | 82.56 | 82.56 | 82.31 | 82.25 | 79.00 | 73.25 | ||||||
| no | 81.08 | 80.94 | 80.64 | 80.14 | 76.94 | 71.56 | ||||||
| pl | 84.00 | 83.97 | 84.06 | 83.39 | 81.28 | 76.83 | ||||||
| pt | 82.42 | 82.50 | 82.00 | 81.58 | 79.36 | 72.97 | ||||||
| ro | 89.36 | 89.39 | 89.22 | 89.14 | 87.03 | 82.06 | ||||||
| ru | 88.97 | 89.17 | 89.08 | 88.69 | 87.00 | 82.17 | ||||||
| sv | 78.06 | 77.86 | 78.00 | 77.47 | 74.83 | 69.53 | ||||||
| th | 81.61 | 81.64 | 81.14 | 80.36 | 76.97 | 68.17 | ||||||
| tr | 81.31 | 81.44 | 81.22 | 80.69 | 77.39 | 71.33 | ||||||
| uk | 88.56 | 88.42 | 88.61 | 87.89 | 85.86 | 80.72 | ||||||
| vi | 86.64 | 86.72 | 86.69 | 85.92 | 83.17 | 77.86 | ||||||
| zh | 78.97 | 79.06 | 78.86 | 78.08 | 74.81 | 67.36 | ||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||
| average | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
| ar | 76.25 | 76.17 | 76.17 | 75.19 | 73.61 | 68.25 | ||||||
| bn | 69.00 | 69.00 | 68.94 | 68.75 | 66.19 | 59.39 | ||||||
| da | 88.53 | 88.39 | 88.25 | 87.58 | 85.69 | 80.25 | ||||||
| de | 92.47 | 92.56 | 92.42 | 92.31 | 90.64 | 87.19 | ||||||
| el | 73.33 | 73.50 | 73.47 | 73.39 | 72.75 | 68.61 | ||||||
| en | 78.58 | 78.61 | 78.36 | 77.86 | 76.44 | 72.58 | ||||||
| es | 86.28 | 86.39 | 86.33 | 85.78 | 84.03 | 80.17 | ||||||
| fi | 84.19 | 84.25 | 84.00 | 83.58 | 82.08 | 77.81 | ||||||
| fr | 90.89 | 90.86 | 90.75 | 90.14 | 88.61 | 84.89 | ||||||
| hi | 61.64 | 61.64 | 61.44 | 61.44 | 58.86 | 52.81 | ||||||
| id | 86.31 | 86.25 | 86.19 | 86.14 | 83.92 | 80.19 | ||||||
| it | 90.17 | 90.08 | 90.17 | 89.83 | 87.78 | 83.75 | ||||||
| ja | 88.50 | 88.64 | 88.64 | 87.92 | 86.22 | 81.64 | ||||||
| ko | 81.42 | 81.42 | 81.31 | 80.53 | 78.31 | 73.03 | ||||||
| nl | 82.47 | 82.44 | 82.36 | 81.94 | 79.97 | 76.00 | ||||||
| no | 83.75 | 83.81 | 83.78 | 83.28 | 80.39 | 75.39 | ||||||
| pl | 84.61 | 84.67 | 84.47 | 83.81 | 82.36 | 78.56 | ||||||
| pt | 83.94 | 83.83 | 83.64 | 83.25 | 81.72 | 77.00 | ||||||
| ro | 91.31 | 91.14 | 91.22 | 91.03 | 89.44 | 85.67 | ||||||
| ru | 90.64 | 90.58 | 90.31 | 90.19 | 88.50 | 84.86 | ||||||
| sv | 78.28 | 78.22 | 78.06 | 77.50 | 75.97 | 71.42 | ||||||
| th | 81.94 | 81.89 | 81.94 | 81.28 | 79.25 | 73.61 | ||||||
| tr | 82.67 | 82.72 | 82.39 | 81.92 | 80.36 | 76.06 | ||||||
| uk | 89.53 | 89.53 | 89.31 | 88.81 | 88.00 | 84.69 | ||||||
| vi | 88.06 | 87.94 | 87.61 | 87.47 | 85.92 | 82.00 | ||||||
| zh | 79.22 | 79.06 | 78.83 | 78.50 | 76.11 | 70.22 | ||||||
| Language - Dimension |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||
| average | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
| de | 85.70 | 85.40 | 84.90 | 84.70 | 81.10 | 79.30 | ||||||
| en | 89.40 | 89.60 | 88.70 | 88.90 | 86.50 | 83.00 | ||||||
| es | 85.90 | 85.90 | 86.00 | 85.70 | 84.00 | 80.80 | ||||||
| fr | 85.10 | 84.70 | 84.80 | 85.30 | 83.30 | 79.10 | ||||||
| it | 85.80 | 85.50 | 86.10 | 85.80 | 83.70 | 81.10 | ||||||
| ko | 82.10 | 81.90 | 82.10 | 80.10 | 77.90 | 71.90 | ||||||
| pl | 86.50 | 86.70 | 86.40 | 86.30 | 84.80 | 80.40 | ||||||
| ru | 81.10 | 81.20 | 80.80 | 81.40 | 77.80 | 74.00 | ||||||
| tr | 83.70 | 84.10 | 83.50 | 83.10 | 80.00 | 75.80 | ||||||
| zh | 83.40 | 83.50 | 82.70 | 81.90 | 78.90 | 73.10 | ||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||
| average | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
| de | 86.20 | 86.40 | 86.10 | 85.90 | 85.20 | 81.50 | ||||||
| en | 90.50 | 90.50 | 91.10 | 90.40 | 90.10 | 86.80 | ||||||
| es | 87.00 | 86.80 | 86.70 | 86.90 | 86.10 | 81.80 | ||||||
| fr | 85.20 | 84.90 | 85.20 | 85.00 | 84.20 | 81.70 | ||||||
| it | 88.00 | 88.00 | 87.70 | 87.60 | 85.70 | 83.20 | ||||||
| ko | 82.10 | 82.20 | 82.40 | 82.80 | 80.30 | 77.50 | ||||||
| pl | 88.50 | 88.70 | 88.60 | 88.30 | 85.90 | 83.60 | ||||||
| ru | 83.10 | 83.00 | 82.70 | 82.80 | 80.80 | 78.10 | ||||||
| tr | 85.60 | 85.60 | 85.90 | 85.90 | 84.50 | 79.80 | ||||||
| zh | 84.10 | 84.10 | 83.80 | 82.80 | 80.90 | 76.20 | ||||||
| Dataset - Dimensions |
| Retrieval - nDCG@10 | ||||||||||||
| Average Retrieval | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
| FiQA2018 | 41.93 | 42.04 | 41.79 | 40.96 | 38.75 | 34.27 | ||||||
| NFCorpus | 32.89 | 33.01 | 32.72 | 32.32 | 30.20 | 25.69 | ||||||
| SciFact | 65.34 | 65.12 | 65.26 | 64.57 | 62.51 | 57.17 | ||||||
| SCIDOCS | 18.90 | 18.85 | 18.84 | 18.46 | 16.98 | 13.84 | ||||||
| CQADupstackRetrieval | 41.20 | 41.17 | 41.06 | 40.47 | 37.97 | 32.42 | ||||||
| Touche2020 | 23.94 | 24.08 | 24.23 | 24.57 | 25.82 | 23.26 | ||||||
| TRECCOVID | 76.73 | 76.54 | 75.98 | 76.01 | 73.10 | 67.47 | ||||||
| FEVER | 84.96 | 84.99 | 84.83 | 84.37 | 81.40 | 69.26 | ||||||
| HotpotQA | 60.14 | 60.13 | 59.71 | 58.06 | 52.14 | 38.97 | ||||||
| DBPedia | 37.44 | 37.42 | 36.98 | 36.12 | 33.21 | 26.34 | ||||||
| NQ | 57.17 | 57.19 | 57.02 | 56.30 | 53.49 | 45.42 | ||||||
| ClimateFEVER | 30.12 | 30.15 | 30.35 | 30.00 | 26.08 | 20.89 | ||||||
| MSMARCO | 37.47 | 37.43 | 37.45 | 37.16 | 35.74 | 32.47 | ||||||
| ArguAna | 43.55 | 43.50 | 43.53 | 42.83 | 41.06 | 37.21 | ||||||
| QuoraRetrieval | 88.14 | 88.13 | 88.06 | 87.88 | 87.10 | 85.19 | ||||||
| STS - Spearman correlation based on cosine similarity | ||||||||||||
| Average STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
| STS12 | 76.72 | 76.72 | 76.85 | 76.97 | 77.13 | 76.44 | ||||||
| STS13 | 79.90 | 79.89 | 79.98 | 80.48 | 80.00 | 78.69 | ||||||
| STS14 | 77.49 | 77.51 | 77.57 | 77.67 | 77.10 | 75.50 | ||||||
| STS15 | 86.42 | 86.43 | 86.46 | 86.58 | 86.21 | 84.58 | ||||||
| STS16 | 85.18 | 85.17 | 85.15 | 85.10 | 84.52 | 84.03 | ||||||
| STS17 | 87.87 | 87.89 | 87.74 | 87.31 | 86.91 | 85.36 | ||||||
| STS22 | 67.07 | 67.04 | 67.01 | 66.95 | 66.66 | 66.71 | ||||||
| BIOSSES | 83.00 | 83.03 | 82.65 | 82.33 | 81.18 | 78.27 | ||||||
| STSBenchmark | 86.87 | 86.68 | 86.86 | 86.72 | 86.26 | 85.14 | ||||||
| SICK-R | 82.38 | 82.38 | 82.36 | 82.27 | 81.87 | 80.89 | ||||||